# Learning State-Invariant Representations of Objects from Image Collections with State, Pose, and Viewpoint Changes

### v1

@@ 6 Conclusion
We present ObjectsWithStateChange, a novel dataset that, in addition to capturing the more commonly used transformations like those related to pose and viewpoint changes, also includes changes in the state of the objects. We evaluate several state-of-the-art methods across eight category-level and object-level classification and retrieval tasks on this dataset. Additionally, we introduce an example mining strategy to learn more discriminative object-level representations for fine-grained tasks. This approach not only boosts performance on our dataset but also on two other multi-view datasets available publicly.
@@
Broader Impact and Future Work: Firstly, we introduce a novel dataset that captures complex state changes of objects that can be used for category-level and fine-grained object-level recognition and retrieval. This could benefit several real-world applications such as automatic checkout systems [38, 39, 9] and robotic systems that deal with object state changes upon interaction. Secondly, it can serve as a real-world test dataset for assessing the robustness of representations learned to state and pose changes, as we show in the supplemental in Appendix B. Lastly, we provide text annotations that can be used to learn invariant multi-modal representations[40, 41] and can be used for generalized zero-shot evaluation, which we aim to pursue in future work.
@@
Appendix A State changes of different objects
@@
A comprehensive list of all possible state changes and other transformations for objects from each category is provided in Table V. Depending on the physical characteristics of the specific object, all the transformations may not be possible for each object, and images are collected if such state changes are possible for the objects of interest.
@@
Appendix B Cross-Dataset Evaluation for Assessing Robustness to State Changes and Other Transformations
To demonstrate this, we use the PiRO model [11] and our method trained on the ObjectPI [7] dataset and test it on the proposed ObjectsWithStateChange dataset and show the results in Table VI. Note that between the two datasets, there are 5 seen and 16 unseen categories and all unseen objects. Even for the seen categories, the object transformations and state changes are not present in the training data and the viewpoints are arbitrary. We observe that our curriculum learning slightly improves performance on object-level tasks.
@@ Appendix B Cross-Dataset Evaluation for Assessing Robustness to State Changes and Other Transformations
TABLE VI: Results of cross-dataset performance evaluation. Model is trained on ObjectPI and tested on the ObjectsWithStateChange dataset with 5 seen and 16 unseen categories and all unseen objects. Even for the seen categories, the object transformations and state changes are unseen, and the viewpoints are arbitrary.
@@
Here we compared the performance of two methods trained on the same ObjectPI dataset and tested on our proposed dataset. In a similar manner, it is also possible to compare the performance of models trained on different datasets to assess how robust and generalizable models trained on different datasets are.
Appendix C Qualitative Fine-grained Retrieval Results
As mentioned earlier in Section 3, our dataset comprises visually similar objects from each of the 21 categories. In this section, we show some qualitative single-image retrieval results of visually similar objects from each category in our dataset in Figs. 8 and 9.
@@ Appendix C Qualitative Fine-grained Retrieval Results
Appendix D Discussion
Broader Impact and Future Research: Firstly, we believe that to the best of our knowledge, we are the first to introduce a dataset that captures complex state changes of objects that can be used for category-level and fine-grained object-level recognition and retrieval tasks. This could benefit several real-world applications such as automatic checkout systems [38, 39, 9], robotic systems, that deal with object state changes upon interaction. Secondly, this dataset can be used for cross-dataset evaluation of methods trained on different datasets and tested for their robustness to transformations captured in our ObjectsWithStateChange dataset. Lastly, we provide text annotations describing the visual characteristics of objects that can be used for training multi-modal models such as [40, 41]. This can be used for adapting such models on object image collections for learning transformation-invariant multi-modal representations and generalized zero-shot evaluation of algorithms, which we aim to pursue in future.
Challenges: The following aspects make our dataset challenging: (a) significant appearance changes due to state changes and other transformations such as arbitrary pose, viewpoint, lighting, and background changes (b) presence of several similar-looking objects with fine-grained differences which makes recognition and retrieval of the correct object-identity challenging especially when there are state changes (c) Additionally, there is class imbalance as some categories such as headphones, telephones, and umbrellas have less number of objects as compared to clothes, decorations, retail objects as objects in the latter categories are more easily available in households. Also, retail objects and decorations have a lot more images than other categories such as cups, bottles, and plates, as many transformations due to pose and state changes are possible in the former than the latter categories. Due to these reasons, the dataset is imbalanced both with respect to the number of objects per category and the number of images per object, which makes it even more challenging. (d) images of the objects are captured in the wild with cluttered and complex backgrounds which facilitates evaluation of algorithms in real-world scenarios.

### v2

@@ 6 Conclusion
We present ObjectsWithStateChange, a novel dataset that, in addition to capturing the more commonly used transformations like those related to pose and viewpoint changes, also includes changes in the state of the objects. We evaluate several state-of-the-art methods across eight category-level and object-level classification and retrieval tasks on this dataset. Additionally, we introduce an example mining strategy to learn more discriminative object-level representations for fine-grained tasks. This approach not only boosts performance on our dataset but also on three other multi-view datasets available publicly.
@@
Broader Impact and Future Work: Firstly, we introduce a novel dataset that captures complex state changes of objects that can be used for category-level and fine-grained object-level recognition and retrieval. This could benefit several real-world applications such as automatic checkout systems [19, 20, 9] and robotic systems that deal with object state changes upon interaction. Secondly, it can serve as a real-world test dataset for assessing the robustness of representations learned to state and pose changes, as we show in the supplement (ref. Appendix C). Lastly, we provide text annotations that can be used to learn invariant multi-modal representations [21, 22, 23] and can be used for generalized zero-shot evaluation and visual question answering, which we aim to pursue in future work.
@@
Appendix A State changes of different objects during data collection
@@
A comprehensive list of all possible state changes and other transformations for objects from each category is provided in Table A5. Depending on the physical characteristics of the specific object, all the transformations may not be possible for each object, and images are collected if such state changes are possible for the objects of interest.
TABLE A5: This table mentions all possible state changes of the objects from 21 categories that are considered during the data collection.
Category State changes and other transformations
bags
place flat on a surface, against wall, hang it on a hook, open and close bag, empty bag and fill it to change shape
books place flat on surfaces, open and close books and place with cover up
bottles open and close lid, fill liquid
bowls
place normally and up-side down, place objects/food inside, open and close lid (if bowl has lid)
clothes put on a hanger, place on floor in a heap, fold properly, hang outside for drying
cups place normally and up-side down, place liquid (if possible)
decorations
place in different stable poses, capture from viewpoints with significant change in appearance, place in complicated backgrounds, and different lighting conditions
(especially ones that have reflective surfaces/have light fixtures), change the arrangement of detachable items (if possible)
headphones
change orientations of microphone, change how the wire is placed, lie flat on table and in an upright position
telephones change receiver position and orientation, place hanging on wall and flat on table
pillows
place on floor and against wall, put force on pillow to deform in several ways, fold (if possible)
plants
move (only if light and not fixed), different lighting conditions (sunlight/cloudy/night), with and without flowers (when applicable)
plates place normally and up-side down, place objects/food on surface
remotes place in different orientations, place in different backgrounds
retail products
place in different poses and backgrounds, change the appearance by removing detachable components, opening cap/lid and placing it beside the object (if possible),
deform objects (if possible)
toys
place in different poses and backgrounds, change the appearance by removing detachable components (if possible), deform objects (if possible)
ties
tie a knot and place normally and in an arbitrary pose, open knot and place normally and in arbitrary pose, roll the tie
towels
fold and place on a surface, fold and place on a hanger, put on a surface in a heap (with different types of deformations), place on arbitrary-shaped surfaces
trolley bags
place on different stable poses (upright on wheels/flat on ground, on different surfaces along length/breadth), with the handle inside/drawn out,
open the bag and place it on the floor and in an upright condition
tumblers
place normally and upside down, fill liquid (if possible), put complicated backgrounds (especially for transparent ones), place in different lighting conditions
(especially for those with reflective surfaces)
umbrellas
open state, closed state with/without tying, fold (if possible), handle inside/drawn out, different backgrounds, folded and hanging
vegetables
raw state, arrange multiple items in different ways, place on different utensils, peel skin (if possible), chop into large chunks (if possible), chop finely (if possible)
Appendix B Time Complexity Analysis
For this analysis, we assume there are $N^{O}$ objects distributed across $N^{C}$ categories in the dataset. The distribution of objects per category is assumed to be uniform, with each category containing an average of $N^{O}_{C}=\frac{N^{O}}{N^{C}}$ objects. Our goal is to examine how the proposed sampling strategy scales with the number of categories and objects. It is important to note that to reduce the complexity of determining neighboring objects in the embedding space, we use a single aggregated embedding for each object extracted from our encoder, regardless of the number of images associated with it. So, the number of images per object is not included in this analysis.
B.1 Analysis of the Proposed Sampling Strategy
At the beginning of each epoch, our algorithm utilizes either of the three distinct sampling strategies:
(S1) For each object, we randomly select another object from the same category. This operation has a time complexity of $\mathcal{O}(1)$. When applied to the entire dataset, the overall time complexity is $\mathcal{O}(N^{O})$.
(S2) For each object, we construct a list of similar objects within the same category and randomly sample an object pair from this list. This requires finding approximate all-nearest-neighbors [24] for all objects in the same category, which incurs a complexity of $\mathcal{O}(N^{O}_{C}\log N^{O}_{C})$ for each category using efficient ANN techniques [18]. Consequently, the total time complexity for the entire dataset is $\mathcal{O}(N^{C}N^{O}_{C}\log N^{O}_{C})=\mathcal{O}(N^{O}\log N^{O}_{C})$.
(S3) This strategy involves partitioning the embedding space using k-means clustering and building an inverted index, which has a time complexity of $\mathcal{O}(KN^{O}I)$, where $K$ representing the number of clusters and $I$ denoting the number of iterations are constants. For each object, we randomly sample another object from the same partition, which has a time complexity of $\mathcal{O}(1)$. Thus, for the entire dataset, the overall time complexity for partitioning and sampling object pairs is $\mathcal{O}(KN^{O}I)\approx\mathcal{O}(N^{O})$, since both $K$ and $I$ are constants independent of $N^{O}$ or $N^{C}$.
Overall, the average time complexity of the algorithm is dominated by the second sampling strategy, yielding $\mathcal{O}(N^{O}\log N^{O}_{C})$ for the entire dataset comprising $N^{O}$ objects. For each object, the time complexity of finding another object for comparison is therefore $\mathcal{O}(\log N^{O}_{C})$, which is sub-linear with respect to the number of objects per category in the dataset.
B.2 Comparison with Other Methods:
In this section, we compare the time complexity per object for different methods. The state-of-the-art method, PiRO [11], randomly selects another object from the same category for each object, resulting in a time complexity of $\mathcal{O}(1)$ per object. In contrast, prior work (PI-TC) [7] compares each object with the nearest multi-view embedding of the object and the nearest proxy embedding for the categories, which incurs a complexity of $\mathcal{O}(N^{C}+N^{O})=\mathcal{O}(N^{O})$ per object, given that $N^{C}<N^{O}$.
Our method employs different sampling strategies across various epochs. Sampling strategies S1 and S3 share the same complexity as PiRO [11], assuming that $K$ and $I$ remain small relative to $N^{O}$. However, S2 exhibits a higher complexity than PiRO due to the additional $logN^{O}_{C}$ term. For most datasets, $N^{O}_{C}$ is relatively small, so the logarithmic term does not significantly impact overall complexity. Nevertheless, for datasets with a large number of objects per category, sampling strategy S2 would have higher complexity than PiRO, although it will still have better complexity than PI-TC [7].
@@ B.2 Comparison with Other Methods:
Appendix C Cross-Dataset Evaluation for Assessing Robustness to State Changes and Other Transformations
@@
TABLE C6: Results of cross-dataset performance evaluation. Model is trained on ObjectPI and tested on the ObjectsWithStateChange dataset with 5 seen and 16 unseen categories and all unseen objects. Even for the seen categories, the object transformations and state changes are unseen, and the viewpoints are arbitrary.
@@
To demonstrate this, we use the PiRO model [11] and our method trained on the ObjectPI [7] dataset and test it on the proposed ObjectsWithStateChange dataset and show the results in Table C6. Note that between the two datasets, there are 5 seen and 16 unseen categories and all unseen objects. Even for the seen categories, the object transformations and state changes are not present in the training data and the viewpoints are arbitrary. Thus, the training data (from ObjectPI) only has variations in pose where the objects are photographed from pre-defined viewpoints, whereas in the test dataset (from ObjectsWithStateChange), we additionally have state changes and complex transformations of objects photographed from arbitrary viewpoints. From Table C6, we observe that our curriculum learning slightly improves performance on object-level tasks.
Here, we compared the performance of two methods trained on the same ObjectPI dataset and tested on our proposed dataset. Using the same setup, it is also possible to compare the performance of models trained on different datasets to assess how robust and generalizable models trained on different datasets are to state changes.
Appendix D Qualitative Fine-grained Retrieval Results
As mentioned earlier in Section 3, our dataset comprises visually similar objects from each of the 21 categories. In this section, we show some qualitative single-image retrieval results of visually similar objects from each category in our dataset in Figs. 11 and 12.
@@ Appendix D Qualitative Fine-grained Retrieval Results
Figure 11: This figure shows single-view object retrieval results of the objects from the first 11 categories. The two columns separated by the thick black line in the middle show results for similar-looking objects within each category. Specifically, in the left column, we show the top-k retrieval results for a SV query image while in the same row of the right column, we show the SV retrieval results for another similar object from the same category. The figure is best viewed when zoomed in. The blue bounding boxes around the images indicate the SV query image while the green and red bounding boxes indicate the correctly and incorrectly retrieved images of the same object-identity as the query image respectively. The results for the other categories are shown on the next page.
Figure 12: This figure shows single-view object retrieval results of the objects from the remaining 10 categories. Following the same format as the previous figure, the two columns separated by the thick black line in the middle show results for similar-looking objects within each category. Specifically, in the left column, we show the top-k retrieval results for a SV query image while in the same row of the right column, we show the SV retrieval results for another similar object from the same category. The figure is best viewed when zoomed in. The blue bounding boxes around the images indicate the SV query image while the green and red bounding boxes indicate the correctly and incorrectly retrieved images of the same object-identity as the query image respectively.
Appendix E Discussion
Broader Impact and Future Research: Firstly, we believe that to the best of our knowledge, we are the first to introduce a dataset that captures complex state changes of objects that can be used for category-level and fine-grained object-level recognition and retrieval tasks. This could benefit several real-world applications such as automatic checkout systems [19, 20, 9], robotic systems, that deal with object state changes upon interaction. Secondly, this dataset can be used for cross-dataset evaluation of methods trained on different datasets and tested for their robustness to state changes and other intricate transformations captured in our ObjectsWithStateChange dataset. Lastly, we provide text annotations describing the visual characteristics of objects that can be used for training multi-modal models such as [21, 22]. This can be used for adapting such models on object image collections for learning transformation-invariant multi-modal representations and generalized zero-shot evaluation of algorithms as well as visual question answering, which we aim to pursue in the future.
Challenges: The following aspects make our dataset challenging: (a) significant appearance changes due to state changes and other transformations such as arbitrary pose, viewpoint, lighting, and background changes (b) presence of several similar-looking objects with fine-grained differences which makes recognition and retrieval of the correct object-identity challenging especially when there are state changes (c) images of the objects are captured in the wild with cluttered and complex backgrounds which facilitates the evaluation of algorithms in real-world scenarios.